# A Survey of Reinforcement Learning For Economics

### 15 Mar 2026 v2

@@ 12.1 Learning Rewards from Preferences
The canonical RLHF framework is built on a formal model of human preference. The observed data consists of tuples $(s,y_{w},y_{l})$, where $s$ is a context, and $y_{w}$ and $y_{l}$ are two outputs, with $y_{w}$ being the “winner” preferred by a human. Assuming preferences follow a latent utility model, the Bradley-Terry model (Bradley and Terry, 1952) gives the probability that $y_{w}$ is preferred: $P(y_{w}\succ y_{l}|s)=\sigma(r_{\theta}(s,y_{w})-r_{\theta}(s,y_{l}))$, where $r_{\theta}:\mathcal{S}\times\mathcal{Y}\to\mathbb{R}$ is a learned reward model parameterized by $\theta$, trained to approximate human preferences (not the ground-truth reward, which is unobserved), and $\sigma(\cdot)$ is the logistic function. This formulation is a binary logit model (Section 2, Equation 1).187187187Iskhakov et al. (2020) discuss the contrasts and synergies between machine learning and structural econometrics, including the shared reliance on logit-based choice models that underlies both RLHF and dynamic discrete choice estimation. The reward model parameters $\theta$ are estimated by minimizing the negative log-likelihood of the observed human choices:
@@
In the LLM setting, the “outputs” $y_{w}$ and $y_{l}$ are token sequences, i.e., trajectories of the autoregressive policy188188188An autoregressive language model generates text token by token: at each step it outputs a distribution over the vocabulary conditioned on all preceding tokens, then samples the next token; the full response is therefore a trajectory in token space and the model acts as a sequential policy over a vocabulary-sized action set., so preferences are over trajectories rather than single actions. The preference loss in Equation (96) is MLE of a choice model where the “alternatives” are trajectory segments and the “choice” is the human-preferred one.
@@ 12.2 The RLHF Pipeline and Direct Optimization
The learned $r_{\theta}$ then serves as a proxy objective for policy optimization. Building on the reward-learning and RL fine-tuning framework of Ziegler et al. (2019) and Stiennon et al. (2020), the canonical three-stage pipeline was formalized by Ouyang et al. (2022). First, a base pretrained model is initialized via supervised fine-tuning (SFT)189189189Pretraining optimizes a language model to predict the next token across a massive text corpus, producing broad linguistic knowledge with no behavioral objective. Supervised fine-tuning (SFT) continues training on a small curated dataset of (prompt, ideal-response) pairs to specialize the model toward the desired task and establish the reference policy $\pi^{SFT}$ from which the KL penalty is measured. on a small dataset of high-quality demonstrations, yielding an initial policy $\pi^{SFT}$. This grounds the model in the desired style and format. The second step is the reward model training as described, using preference data generated from this $\pi^{SFT}$.
In the third step, the SFT policy $\pi_{\phi}$ is fine-tuned via PPO (Section 5.5) to maximize the frozen reward model $r_{\theta}$, with a KL-divergence penalty preventing the policy from drifting into regions where $r_{\theta}$ is unreliable. The objective is
@@ 12.3 Recent Developments
While RLHF and its successors have proven effective at aligning model behavior with human preferences, several open issues remain. The framework does not aim to solve the identification problem in the strict econometric sense; the learned reward model $r_{\theta}$ (whether explicit or implicit in DPO) is a proxy for preference, not necessarily the uniquely identified “true” utility function required for welfare analysis.191191191The identification problem in preference learning mirrors that in discrete choice. The reward function is identified only up to an additive constant and requires a location normalization. Furthermore, the notion of “human feedback” is a significant simplification, as the preferences being optimized are those of a small, non-representative group of paid labelers, raising important questions about whose values are being embedded in these systems. Finally, the fine-tuning process can lead to degraded performance on standard academic benchmarks, a phenomenon called the alignment tax (Ouyang et al., 2022).192192192The alignment tax, as described by Ouyang et al. (2022), refers to performance regressions on public NLP benchmarks (SQuAD, DROP, HellaSwag) that result from RLHF fine-tuning. Mitigating these challenges while scaling alignment beyond the limits of human data collection remains a key frontier for the field.
@@ 12.4 Simulation Study: Preference Learning in Job Search
Six methods are compared across $K\in\{25,50,100,200,500,1{,}000,2{,}000,5{,}000\}$, averaged over 30 seeds. The neural network RLHF reward model is a two-layer MLP trained by Bradley-Terry MLE on the $K$ segment pairs; per-transition rewards are discount-weighted and summed to obtain a segment score, and the resulting 112-state reward table is solved by value iteration.193193193The neural network has 4 inputs (normalized log-wage, normalized amenity, employment indicator, action), 32 hidden units per layer, and $\sim$1,200 parameters. The logistic loss from Equation (96) is applied to discount-weighted segment reward sums. The correctly specified structural model parameterizes utility as $\hat{u}=\hat{\alpha}\log(w)+(1-\hat{\alpha})z$, estimating the single parameter $\hat{\alpha}$ via Bradley-Terry MLE, then solves the MDP. The misspecified model uses $\hat{u}=\hat{\alpha}\log(w)+(1-\hat{\alpha})\bar{z}$, where $\bar{z}=3$ is the mean amenity level; this model ignores amenity variation across jobs, treating all amenities as identical. DPO trains a tabular softmax policy directly from trajectory comparisons, bypassing reward modeling entirely.194194194DPO uses 112 logit parameters $\phi_{s}$, one per state, trained via the DPO loss (Equation 100) with Adam optimization over a sweep of $\lambda_{KL}\in\{0.01,0.05,0.1,0.5,1.0,5.0\}$, selecting the $\lambda_{KL}$ that minimizes training loss. The reference policy is uniform: $\pi^{SFT}(a|s)=0.5$. To match the LLM setup where both completions condition on the same prompt, DPO comparison pairs start from the same initial state. Tabular Q-learning (10,00010{,}000 episodes, $\varepsilon=0.15$, learning rate $0.1$) and exact DP provide scalar-reward baselines. All four preference methods receive identical comparison data per seed; DPO receives a same-state variant generated from the same seed.
Table 23: Sample complexity of preference-based policy learning on a McCall-style job search MDP with compensating differentials (30 seeds, $L=15$). Six methods spanning four preference-learning approaches (structural, neural network, DPO) and two scalar-reward baselines (DP, Q-learning).
Method $K$ $V^{\pi}(s_{0})$ % of DP
DP Oracle $74.13$ $100.0$
Q-Learning $73.60\pm 0.05$ $99.3$
NN RLHF $25$ $73.05\pm 0.13$ $98.5$
DPO $25$ $66.39\pm 0.34$ $89.6$
Correct $25$ $74.03\pm 0.03$ $99.9$
Misspecified $25$ $67.34\pm 0.38$ $90.8$
NN RLHF $50$ $73.31\pm 0.08$ $98.9$
DPO $50$ $68.02\pm 0.23$ $91.8$
Correct $50$ $74.07\pm 0.02$ $99.9$
Misspecified $50$ $67.12\pm 0.33$ $90.6$
NN RLHF $100$ $73.60\pm 0.06$ $99.3$
DPO $100$ $69.25\pm 0.18$ $93.4$
Correct $100$ $74.10\pm 0.01$ $100.0$
Misspecified $100$ $67.24\pm 0.28$ $90.7$
NN RLHF $200$ $73.85\pm 0.05$ $99.6$
DPO $200$ $70.16\pm 0.16$ $94.6$
Correct $200$ $74.12\pm 0.00$ $100.0$
Misspecified $200$ $67.54\pm 0.17$ $91.1$
NN RLHF $500$ $73.99\pm 0.02$ $99.8$
DPO $500$ $70.60\pm 0.10$ $95.2$
Correct $500$ $74.13\pm 0.00$ $100.0$
Misspecified $500$ $67.55\pm 0.15$ $91.1$
NN RLHF $1000$ $73.95\pm 0.03$ $99.8$
DPO $1000$ $70.59\pm 0.09$ $95.2$
Correct $1000$ $74.12\pm 0.00$ $100.0$
Misspecified $1000$ $67.52\pm 0.12$ $91.1$
NN RLHF $2000$ $74.04\pm 0.02$ $99.9$
DPO $2000$ $70.41\pm 0.07$ $95.0$
Correct $2000$ $74.13\pm 0.00$ $100.0$
Misspecified $2000$ $67.41\pm 0.11$ $90.9$
NN RLHF $5000$ $74.09\pm 0.01$ $99.9$
DPO $5000$ $70.14\pm 0.07$ $94.6$
Correct $5000$ $74.13\pm 0.00$ $100.0$
Misspecified $5000$ $67.46\pm 0.11$ $91.0$
@@
Figure 23: Policy value $V^{\pi}(s_{0})$ versus number of preference comparisons $K$ for all six methods.
@@
Table 23 and Figure 23 report the main results. Three findings stand out. First, the correctly specified structural model dominates: even at $K=25$ it achieves 99.9% of DP-optimal, illustrating the power of correct specification when the model has a single free parameter. Second, the neural network converges more slowly but reaches 99.9% by K=5, [^0] K=5{,}000, reflecting the higher sample complexity of a flexible model relative to a one-parameter structural specification. Third, DPO plateaus at approximately 95% by $K=500$ and does not improve further.195195195DPO learns only from $(s,a)$ pairs visited in training trajectories generated by the random behavioral policy. It cannot propagate value to undervisited states the way value iteration does after learning a reward model, so states poorly covered by the behavioral policy remain suboptimal regardless of $K$. This plateau is qualitatively different from the gridworld failure ($118\%$): DPO recovers a reasonable policy, but the gap does not close with additional data.196196196DPO fails catastrophically in gridworld because transitions are stochastic (10% slip probability) and rewards are transition-dependent. The same $(s,a)$ pair yields different rewards depending on whether the agent slipped, so the DPO loss conflates policy quality with transition luck. In the job search model, accept/reject deterministically changes employment status, and only the 5% layoff probability introduces stochastic transitions. The misspecified constant-amenity model plateaus at 91% regardless of $K$, because additional preference data cannot correct the omitted variable.
@@
Table 24 provides state-level diagnostics at K=5, [^0] K=5{,}000. The structural model and neural network achieve near-perfect policy agreement with $\pi^{*}$ (100% and 96% respectively). DPO agrees on only 57% of states with value-function correlation 0.78, systematically underselecting on both wage and amenity dimensions.197197197DPO’s mean accepted amenity of 3.4 parallels the misspecified structural model’s 3.0, though the mechanisms differ: the misspecified model ignores amenity variation by construction, while DPO underweights amenities because the random behavioral policy underrepresents high-amenity employed states in the training data. The misspecified model agrees on 50%, with disagreements concentrated where amenity variation matters.198198198An online versus offline ablation for the neural network at K=1, [^0] K=1{,}000 (20 seeds) shows comparable performance: online $73.92\pm 0.05$, offline $73.99\pm 0.03$ ($p=0.09$). The random behavioral policy already provides diverse career trajectories covering the full wage-amenity space.
Table 25: Segment length ablation at K=2, [^0] K=2{,}000 (20 seeds). Policy value $V^{\pi}(s_{0})$ for the neural network and DPO across segment lengths $L\in\{1,3,5,10,15,25\}$.
$L$ NN RLHF $V^{\pi}(s_{0})$ % of DP DPO $V^{\pi}(s_{0})$ % of DP
11 $56.62\pm 2.28$ $76.4$ $49.66\pm 0.25$ $67.0$
33 $73.89\pm 0.04$ $99.7$ $69.31\pm 0.11$ $93.5$
55 $74.01\pm 0.02$ $99.8$ $70.04\pm 0.09$ $94.5$
$10$ $74.02\pm 0.03$ $99.9$ $70.29\pm 0.08$ $94.8$
$15$ $74.05\pm 0.02$ $99.9$ $70.57\pm 0.07$ $95.2$
$30$ $74.05\pm 0.02$ $99.9$ $70.53\pm 0.13$ $95.1$
Figure 24: Policy value $V^{\pi}(s_{0})$ versus segment length $L$ at K=2, [^0] K=2{,}000 for the neural network and DPO.
Table 25 and Figure 24 report the segment length ablation at K=2, [^0] K=2{,}000. At $L=1$, both methods perform poorly because single-step comparisons carry minimal information about long-run value. The neural network recovers rapidly (by $L=3$) because the reward model aggregates per-transition estimates over longer segments and value iteration propagates them through the full transition structure; DPO improves monotonically but plateaus at its $\sim$95% ceiling, as it must learn the policy directly from trajectory comparisons without access to the transition model.
Two-stage RLHF separates preference estimation (a static econometric problem) from dynamic programming (which exploits the known transition model); DPO conflates the two and forfeits the transition structure that economists typically have access to.199199199This advantage is specific to settings where the transition model is known or estimable; in domains without a tractable transition model, DPO’s single-stage approach avoids compounding errors from reward model estimation.
@@ 13 Reinforcement Learning and Causal Inference
Reinforcement learning algorithms solve Markov decision processes by estimating value functions or optimizing policies from sampled transitions. Two assumptions underlie the standard formulation. First, the agent’s action at time $t$ is determined solely by the observed state $s_{t}$ and the agent’s policy $\pi(a\mid s_{t})$; there is no unobserved variable (confounder) simultaneously influencing both the action and the reward or state transition.200200200See Section 2 for the terminological mapping between “outcome” in causal inference and the corresponding RL quantities. Throughout this chapter, I reserve “outcome” for its causal inference meaning and use the specific RL quantity (reward, state, return, value) elsewhere. Second, the observed state $s_{t}$ is sufficient for prediction, so that conditioning on $s_{t}$ renders future states independent of past history. Both assumptions are the Markov property restated in causal language. Both fail routinely in economic applications.
@@ 13.2 The Confounded MDP
When unobserved confounders influence both the behavioral policy and the transitions or rewards, the MDP is confounded. This formalization, developed by Zhang and Bareinboim (2019), Zhang and Bareinboim (2020), and Kallus and Zhou (2020), provides the foundation for causal reasoning in sequential decision problems. The key tool is Pearl’s do-operator. The interventional distribution $P(Y\mid\operatorname{do}(X=x))$ is the distribution that arises when $X$ is set externally rather than observed passively, severing all incoming causal influences on $X$ while leaving the remaining data-generating process intact.201201201The do-operator is formalized within the structural causal model (SCM) framework of Pearl (2009). An SCM specifies endogenous variables $\mathbf{V}$, exogenous variables $\mathbf{U}$, structural equations $V_{i}=f_{i}(\text{pa}(V_{i}),U_{i})$, and a distribution $P(\mathbf{U})$. The intervention $\operatorname{do}(X=x)$ replaces the structural equation for $X$ with a constant, producing the interventional distribution. See Pearl (2009) for the complete framework, including the causal hierarchy (association, intervention, counterfactual) and general identification theory.
@@
The importance-sampling estimator $\hat{V}^{\pi}_{\text{IS}}=\frac{1}{N}\sum_{i=1}^{N}\prod_{t=0}^{T}\frac{\pi(a_{t}^{(i)}\mid s_{t}^{(i)})}{\mu(a_{t}^{(i)}\mid s_{t}^{(i)})}G^{(i)}$, where $G^{(i)}=\sum_{t=0}^{T}\gamma^{t}R_{t}^{(i)}$ is the discounted return of trajectory $i$ and $T$ is the trajectory length, is also biased because the propensity $\mu(a\mid s)$ is not the true behavioral propensity $\mu(a\mid s,u)$.202202202This is the sequential analogue of the omitted variable bias in linear regression. In the static case, regressing $Y$ on $X$ without controlling for a confounder $U$ yields a biased coefficient. In the sequential case, the bias propagates through the Bellman recursion and can amplify over the horizon.
@@
Suppose a set of observed variables $\mathbf{Z}_{t}$ satisfies the backdoor criterion relative to $(A_{t},S_{t+1})$ in the causal graph of the confounded MDP, meaning that $\mathbf{Z}_{t}$ blocks all backdoor paths from $A_{t}$ to $S_{t+1}$ and no element of $\mathbf{Z}_{t}$ is a descendant of $A_{t}$.203203203A backdoor path from $A_{t}$ to $S_{t+1}$ is any path in the causal graph that begins with an arrow into $A_{t}$ (i.e., a non-causal path). In the confounded MDP, $A_{t}\leftarrow U_{t}\rightarrow S_{t+1}$ is a backdoor path: $U_{t}$ causes both $A_{t}$ and $S_{t+1}$, creating a spurious association. Blocking all such paths by conditioning on appropriate variables eliminates the confounding bias. See Pearl (2009), Chapter 3. Then the interventional transition probability is identified:

### 17 Mar 2026 v3

@@ 11.5 Learning Rewards from Preferences
The canonical RLHF framework is built on a formal model of human preference. The observed data consists of tuples $(s,y_{w},y_{l})$, where $s$ is a context, and $y_{w}$ and $y_{l}$ are two outputs, with $y_{w}$ being the “winner” preferred by a human. Assuming preferences follow a latent utility model, the Bradley-Terry model (Bradley and Terry, 1952) gives the probability that $y_{w}$ is preferred: $P(y_{w}\succ y_{l}|s)=\sigma(r_{\theta}(s,y_{w})-r_{\theta}(s,y_{l}))$, where $r_{\theta}:\mathcal{S}\times\mathcal{Y}\to\mathbb{R}$ is a learned reward model parameterized by $\theta$, trained to approximate human preferences (not the ground-truth reward, which is unobserved), and $\sigma(\cdot)$ is the logistic function. This formulation is a binary logit model (Section 2, Equation 1).203203203Iskhakov et al. (2020) discuss the contrasts and synergies between machine learning and structural econometrics, including the shared reliance on logit-based choice models that underlies both RLHF and dynamic discrete choice estimation. The reward model parameters $\theta$ are estimated by minimizing the negative log-likelihood of the observed human choices:
@@
In the LLM setting, the “outputs” $y_{w}$ and $y_{l}$ are token sequences, i.e., trajectories of the autoregressive policy204204204An autoregressive language model generates text token by token: at each step it outputs a distribution over the vocabulary conditioned on all preceding tokens, then samples the next token; the full response is therefore a trajectory in token space and the model acts as a sequential policy over a vocabulary-sized action set., so preferences are over trajectories rather than single actions. The preference loss in Equation (104) is MLE of a choice model where the “alternatives” are trajectory segments and the “choice” is the human-preferred one.
@@ 11.6 The RLHF Pipeline and Direct Optimization
The learned $r_{\theta}$ then serves as a proxy objective for policy optimization. Building on the reward-learning and RL fine-tuning framework of Ziegler et al. (2019) and Stiennon et al. (2020), the canonical three-stage pipeline was formalized by Ouyang et al. (2022). First, a base pretrained model is initialized via supervised fine-tuning (SFT)205205205Pretraining optimizes a language model to predict the next token across a massive text corpus, producing broad linguistic knowledge with no behavioral objective. Supervised fine-tuning (SFT) continues training on a small curated dataset of (prompt, ideal-response) pairs to specialize the model toward the desired task and establish the reference policy $\pi^{SFT}$ from which the KL penalty is measured. on a small dataset of high-quality demonstrations, yielding an initial policy $\pi^{SFT}$. This grounds the model in the desired style and format. The second step is the reward model training as described, using preference data generated from this $\pi^{SFT}$.
In the third step, the SFT policy $\pi_{\phi}$ is fine-tuned via PPO (Section 5.5) to maximize the frozen reward model $r_{\theta}$,206206206After fine-tuning concludes, the resulting LLM is deployed with frozen weights. Each user interaction is a forward pass in the execution phase (Section 2); the model does not update its parameters from conversations. Periodic retraining on new preference data constitutes a separate training phase. with a KL-divergence penalty preventing the policy from drifting into regions where $r_{\theta}$ is unreliable. The objective is
@@ 11.7 Recent Developments
While RLHF and its successors have proven effective at aligning model behavior with human preferences, several open issues remain. The framework does not aim to solve the identification problem in the strict econometric sense; the learned reward model $r_{\theta}$ (whether explicit or implicit in DPO) is a proxy for preference, not necessarily the uniquely identified “true” utility function required for welfare analysis.208208208The identification problem in preference learning mirrors that in discrete choice. The reward function is identified only up to an additive constant and requires a location normalization. Furthermore, the notion of “human feedback” is a significant simplification, as the preferences being optimized are those of a small, non-representative group of paid labelers, raising important questions about whose values are being embedded in these systems. Finally, the fine-tuning process can lead to degraded performance on standard academic benchmarks, a phenomenon called the alignment tax (Ouyang et al., 2022).209209209The alignment tax, as described by Ouyang et al. (2022), refers to performance regressions on public NLP benchmarks (SQuAD, DROP, HellaSwag) that result from RLHF fine-tuning. Mitigating these challenges while scaling alignment beyond the limits of human data collection remains a key frontier for the field.
@@ 11.8 Simulation Study: Preference Learning in Job Search
Six methods are compared across $K\in\{25,50,100,200,500,1{,}000,2{,}000,5{,}000\}$, averaged over 30 seeds. The neural network RLHF reward model is a two-layer MLP trained by Bradley-Terry MLE on the $K$ segment pairs; per-transition rewards are discount-weighted and summed to obtain a segment score, and the resulting 112-state reward table is solved by value iteration.210210210The neural network has 4 inputs (normalized log-wage, normalized amenity, employment indicator, action), 32 hidden units per layer, and $\sim$1,200 parameters. The logistic loss from Equation (104) is applied to discount-weighted segment reward sums. The correctly specified structural model parameterizes utility as $\hat{u}=\hat{\alpha}\log(w)+(1-\hat{\alpha})z$, estimating the single parameter $\hat{\alpha}$ via Bradley-Terry MLE, then solves the MDP. The misspecified model uses $\hat{u}=\hat{\alpha}\log(w)+(1-\hat{\alpha})\bar{z}$, where $\bar{z}=3$ is the mean amenity level; this model ignores amenity variation across jobs, treating all amenities as identical. DPO trains a tabular softmax policy directly from trajectory comparisons, bypassing reward modeling entirely.211211211DPO uses 112 logit parameters $\phi_{s}$, one per state, trained via the DPO loss (Equation 108) with Adam optimization over a sweep of $\lambda_{KL}\in\{0.01,0.05,0.1,0.5,1.0,5.0\}$, selecting the $\lambda_{KL}$ that minimizes training loss. The reference policy is uniform: $\pi^{SFT}(a|s)=0.5$. To match the LLM setup where both completions condition on the same prompt, DPO comparison pairs start from the same initial state. Tabular Q-learning (10,00010{,}000 episodes, $\varepsilon=0.15$, learning rate $0.1$) and exact DP provide scalar-reward baselines. All four preference methods receive identical comparison data per seed; DPO receives a same-state variant generated from the same seed.
@@
Figure 23: Policy value $V^{\pi}(s_{0})$ versus number of preference comparisons $K$ for all six methods (30 seeds, $L=15$). The right axis shows the percentage of DP-optimal value.
@@
Figure 23 reports the main results. Three findings stand out. First, the correctly specified structural model dominates: even at $K=25$ it achieves 99.9% of DP-optimal, illustrating the power of correct specification when the model has a single free parameter. Second, the neural network converges more slowly but reaches 99.9% by K=5, [^0] K=5{,}000, reflecting the higher sample complexity of a flexible model relative to a one-parameter structural specification. Third, DPO plateaus at approximately 95% by $K=500$ and does not improve further.212212212DPO learns only from $(s,a)$ pairs visited in training trajectories generated by the random behavioral policy. It cannot propagate value to undervisited states the way value iteration does after learning a reward model, so states poorly covered by the behavioral policy remain suboptimal regardless of $K$. This plateau is qualitatively different from the gridworld failure ($118\%$): DPO recovers a reasonable policy, but the gap does not close with additional data.213213213DPO fails catastrophically in gridworld because transitions are stochastic (10% slip probability) and rewards are transition-dependent. The same $(s,a)$ pair yields different rewards depending on whether the agent slipped, so the DPO loss conflates policy quality with transition luck. In the job search model, accept/reject deterministically changes employment status, and only the 5% layoff probability introduces stochastic transitions. The misspecified constant-amenity model plateaus at 91% regardless of $K$, because additional preference data cannot correct the omitted variable.
@@
Table 21 provides state-level diagnostics at K=5, [^0] K=5{,}000. The structural model and neural network achieve near-perfect policy agreement with $\pi^{*}$ (100% and 96% respectively). DPO agrees on only 57% of states with value-function correlation 0.78, systematically underselecting on both wage and amenity dimensions.214214214DPO’s mean accepted amenity of 3.4 parallels the misspecified structural model’s 3.0, though the mechanisms differ: the misspecified model ignores amenity variation by construction, while DPO underweights amenities because the random behavioral policy underrepresents high-amenity employed states in the training data. The misspecified model agrees on 50%, with disagreements concentrated where amenity variation matters.215215215An online versus offline ablation for the neural network at K=1, [^0] K=1{,}000 (20 seeds) shows comparable performance: online $73.92\pm 0.05$, offline $73.99\pm 0.03$ ($p=0.09$). The random behavioral policy already provides diverse career trajectories covering the full wage-amenity space.
Figure 24: Policy value $V^{\pi}(s_{0})$ versus segment length $L$ at K=2, [^0] K=2{,}000 for the neural network and DPO (20 seeds). The right axis shows the percentage of DP-optimal value.
Figure 24 reports the segment length ablation at K=2, [^0] K=2{,}000. At $L=1$, both methods perform poorly because single-step comparisons carry minimal information about long-run value. The neural network recovers rapidly (by $L=3$) because the reward model aggregates per-transition estimates over longer segments and value iteration propagates them through the full transition structure; DPO improves monotonically but plateaus at its $\sim$95% ceiling, as it must learn the policy directly from trajectory comparisons without access to the transition model.
Two-stage RLHF separates preference estimation (a static econometric problem) from dynamic programming (which exploits the known transition model); DPO conflates the two and forfeits the transition structure that economists typically have access to.216216216This advantage is specific to settings where the transition model is known or estimable; in domains without a tractable transition model, DPO’s single-stage approach avoids compounding errors from reward model estimation.
@@ 12 Reinforcement Learning and Causal Inference
Reinforcement learning algorithms solve Markov decision processes by estimating value functions or optimizing policies from sampled transitions. Two assumptions underlie the standard formulation. First, the agent’s action at time $t$ is determined solely by the observed state $s_{t}$ and the agent’s policy $\pi(a\mid s_{t})$; there is no unobserved variable (confounder) simultaneously influencing both the action and the reward or state transition.217217217See Section 2 for the terminological mapping between “outcome” in causal inference and the corresponding RL quantities. Throughout this chapter, I reserve “outcome” for its causal inference meaning and use the specific RL quantity (reward, state, return, value) elsewhere. Second, the observed state $s_{t}$ is sufficient for prediction, so that conditioning on $s_{t}$ renders future states independent of past history. Both assumptions are the Markov property restated in causal language. Both fail routinely in economic applications.
@@ 12.2 The Confounded MDP
When unobserved confounders influence both the behavioral policy and the transitions or rewards, the MDP is confounded. This formalization, developed by Zhang and Bareinboim (2019), Zhang and Bareinboim (2020), and Kallus and Zhou (2020), provides the foundation for causal reasoning in sequential decision problems. The key tool is Pearl’s do-operator. The interventional distribution $P(Y\mid\operatorname{do}(X=x))$ is the distribution that arises when $X$ is set externally rather than observed passively, severing all incoming causal influences on $X$ while leaving the remaining data-generating process intact.218218218The do-operator is formalized within the structural causal model (SCM) framework of Pearl (2009). An SCM specifies endogenous variables $\mathbf{V}$, exogenous variables $\mathbf{U}$, structural equations $V_{i}=f_{i}(\text{pa}(V_{i}),U_{i})$, and a distribution $P(\mathbf{U})$. The intervention $\operatorname{do}(X=x)$ replaces the structural equation for $X$ with a constant, producing the interventional distribution. See Pearl (2009) for the complete framework, including the causal hierarchy (association, intervention, counterfactual) and general identification theory.
@@
The importance-sampling estimator $\hat{V}^{\pi}_{\text{IS}}=\frac{1}{N}\sum_{i=1}^{N}\prod_{t=0}^{T}\frac{\pi(a_{t}^{(i)}\mid s_{t}^{(i)})}{\mu(a_{t}^{(i)}\mid s_{t}^{(i)})}G^{(i)}$, where $G^{(i)}=\sum_{t=0}^{T}\gamma^{t}R_{t}^{(i)}$ is the discounted return of trajectory $i$ and $T$ is the trajectory length, is also biased because the propensity $\mu(a\mid s)$ is not the true behavioral propensity $\mu(a\mid s,u)$.219219219This is the sequential analogue of the omitted variable bias in linear regression. In the static case, regressing $Y$ on $X$ without controlling for a confounder $U$ yields a biased coefficient. In the sequential case, the bias propagates through the Bellman recursion and can amplify over the horizon.
@@
Suppose a set of observed variables $\mathbf{Z}_{t}$ satisfies the backdoor criterion relative to $(A_{t},S_{t+1})$ in the causal graph of the confounded MDP, meaning that $\mathbf{Z}_{t}$ blocks all backdoor paths from $A_{t}$ to $S_{t+1}$ and no element of $\mathbf{Z}_{t}$ is a descendant of $A_{t}$.220220220A backdoor path from $A_{t}$ to $S_{t+1}$ is any path in the causal graph that begins with an arrow into $A_{t}$ (i.e., a non-causal path). In the confounded MDP, $A_{t}\leftarrow U_{t}\rightarrow S_{t+1}$ is a backdoor path: $U_{t}$ causes both $A_{t}$ and $S_{t+1}$, creating a spurious association. Blocking all such paths by conditioning on appropriate variables eliminates the confounding bias. See Pearl (2009), Chapter 3. Then the interventional transition probability is identified: